4. (15) A saturated solution of phenol, C6H5OH, a very weak acid with pKa=10.00, has pH=4.90. What is the molarity of phenol in this solution?
Answer #1

use:

pKa = -log Ka

10.0 = -log Ka

Ka = 1*10^-10

Let the concentration of C6H5OH be c

use:

pH = -log [H+]

4.9 = -log [H+]

[H+] = 1.259*10^-5 M

C6H5OH dissociates as:

C6H5OH -----> H+ + C6H5O-

c 0 0

c-x x x

Ka = [H+][C6H5O-]/[C6H5OH]

Ka = x*x/(c-x)

1*10^-10 = 1.259*10^-5*1.259*10^-5/(c-1.259*10^-5)

c-1.259*10^-5 = 1.585

c=1.585 M

Answer: 1.58 M